# An LLM-based Agent for Reliable Docker Environment Configuration

### v2

@@ Abstract
Environment configuration is a critical yet time-consuming step in software development, especially when dealing with unfamiliar code repositories. While Large Language Models (LLMs) demonstrate the potential to accomplish software engineering tasks, existing methods for environment configuration often rely on manual efforts or fragile scripts, leading to inefficiencies and unreliable outcomes. We introduce Repo2Run, the first LLM-based agent designed to fully automate environment configuration and generate executable Dockerfiles for arbitrary Python repositories. We address two major challenges: (1) enabling the LLM agent to configure environments within isolated Docker containers, and (2) ensuring the successful configuration process is recorded and accurately transferred to a Dockerfile without error. To achieve this, we propose atomic configuration synthesis, featuring a dual-environment architecture (internal and external environment) with a rollback mechanism to prevent environment “pollution” from failed commands, guaranteeing atomic execution (execute fully or not at all) and a Dockerfile generator to transfer successful configuration steps into runnable Dockerfiles. We evaluate Repo2Run on our proposed benchmark of 420 recent Python repositories with unit tests, where it achieves an 86.0% success rate, outperforming the best baseline by 63.9%. Repo2Run is available at https://github.com/bytedance/Repo2Run.
@@ 1 Introduction
Figure 1: The illustration of a command executing failed and “polluting” the environment.
Figure 2: Examples of the failed command causing “polluted” environment.
Configuring the environment is typically the first step for developers to run and test the code. Meanwhile, Large Language Models (LLMs) are increasingly being integrated into tools such as chatbots and coding assistants, showcasing their potential to accomplish various software engineering tasks [32, 12, 4]. As a result, the research community explores how LLMs can be leveraged to assist with more complex real-world tasks in software development. Furthermore, LLM-based agents [13, 5, 1] are also increasingly being utilized in software engineering. With the advancement of LLM-based agents and their applications in automated programming, many techniques now require running tests in the runtime environment to acquire more comprehensive information. However, previous approaches often rely on manually configuring environments [14], and writing individual configuration scripts for each task, which is both time-consuming and labor-intensive.
We conduct an empirical study with eight professional developers from Internet companies, all with an average of seven years of development experience. The results indicate that when dealing with unfamiliar code repositories, five out of eight developers spend a great amount of time configuring the environment. Moreover, there is a strong willingness among them to use automated tools to alleviate the burden of environment setup. Details of the empirical study are presented in Appendix A.
In response to the growing complexity and time required for manual environment configuration, containerization tools like Docker have become a good solution. Docker allows developers to create isolated and consistent environments across different systems, thus eliminating the common “it works on my machine” problem [23]. Dockerfiles are commonly used to configure Docker environments, which include two parts: the base image and the configuration process. The base image is a pre-packaged environment (e.g., python:3.10) that provides a standardized, isolated environment for development without the need for local deployment. The configuration process involves additional steps performed on top of the base image.
Inspired by the idea of the Docker container, we design our LLM agent for environment configuration, Repo2Run, with a fundamental concept: utilize an LLM agent to configure the environment within an isolated Docker container. If the LLM successfully configures the environment, we only need to record the configuration steps and transfer them into a Dockerfile. During the environment configuration process, we mainly face the following two major challenges:
Ensure the LLM agent configures the environment within a Docker container. One major challenge is ensuring that the LLM agent can accurately configure the environment inside the Docker container. This involves selecting an appropriate base image, installing dependencies, resolving conflicts, and running tests, etc. If we do not provide the LLM agent with effective tools, the LLM agent will find it difficult to handle the various complex issues encountered during configuration.
Ensure the successful configuration process is recorded and accurately transferred to a runnable Dockerfile. Another major challenge is maintaining an accurate and comprehensive record of the environment configuration process. As illustrated in Figure 1, a failed executed command “commandA” may cause irreversible ”pollution”. Although “commandA” fails to run, it still causes ”pollution” to the original environment, potentially leading to unexpected changes to packages, files, or directories in the system. To accurately reproduce such system changes, we need to add the “RUN commandA” statement in the Dockerfile. However, this will cause the Dockerfile to build fail.
Figure 2 shows two practical examples. As shown in Figure 2 (a), when we run “pip install tensorflow-gpu”, the command fails to execute because the package tensorflow-gpu has been deprecated. However, it still introduces the “python-version” package which is not present in the original environment. In Figure 2 (b), when we run the “rm -rf /path/to/logs” command to delete the directory, some files fail to be deleted due to a lack of permission. However, the files that have already been deleted (e.g., log1.txt and log2.txt) are not restored. Such “pollution” caused by failed commands may make the environment enter an uncertain state. This state cannot be reached by directly adding commands to the Dockerfile, as doing so would cause the Dockerfile build process to fail.
To address the aforementioned two challenges, we propose Repo2Run based on atomic configuration synthesis, which consists of a dual-environment architecture and a Dockerfile generator. To the best of our knowledge, this is the first work that leverages the LLM agent for automated Dockerfile generation and environment configuration in a completely automated manner.
The dual-environment architecture comprises both an internal and an external environment. The internal environment serves as a Docker container sandbox, allowing the LLM agent to execute various commands. The external environment assists in configuring the environment in the internal environment. It facilitates the action-observation interactions of the LLM agent and supports operations such as changing the base image of the internal environment. In addition, we design a rollback mechanism that performs a rollback when a command fails, preventing the system from entering an uncertain state. This ensures that we can execute commands “atomically”, meaning each command either executes fully or not at all.
The Dockerfile generator designs a set of rules to transfer the successfully executed commands in the internal environment into statements in the Dockerfile without errors. Combined with the rollback mechanism, we finally ensure that all generated Dockerfiles are successfully built.
In summary, the contributions of this paper are as follows:
$\bullet$ We propose Repo2Run, the first LLM-based agent designed to automate environment configuration and generate Dockerfiles for arbitrary Python repositories, enabling the execution of tests within these repositories.
$\bullet$ We propose atomic configuration synthesis, which includes a dual-environment architecture and a Dockerfile generator, ensuring that all Dockerfiles generated by Repo2Run are runnable.
$\bullet$ We create a benchmark to evaluate the performance of Repo2Run by crawling 420 popular Python repositories with unit tests from GitHub released within 2024. Repo2Run successfully configures the environment for 361 of these repositories, achieving a success rate of 86.0%, which is 63.9% higher than the best baseline.
2 Environment configuration
In this part, we formulate the task of the environment configuration. It involves identifying an appropriate base image ($B$) and a configuration process ($\Gamma$). The objective is to ensure that the final system state ($S_{f}$) successfully runs all the unit tests in the repository.
@@ 2.1 State Transition
Environment State ($S$): The environmental state ($S$) represents the current state of the computer system, which encompasses all variables, files, cache, etc. Any state $S\in\mathcal{S}$, where ($\mathcal{S}$) denotes the set of all possible environment states.
Command ($C$): The command $C$ represents an individual instruction or action that can be executed in the environment through interfaces such as bash, thus directly changing the system state.
State Transition Function ($\delta$): The state transition function ($\delta$) defines the process through which the system transitions from one state to another state upon execution of a given command, i.e.,
$\delta:\mathcal{S}\times\mathcal{C}\rightarrow\mathcal{S},\quad\delta(S,C)=S^{\prime}$ (1)
Here, $S^{\prime}$ is the new state after execution of command $C$.
@@ 2.2 Base image
Empty initial state ($S_{\emptyset}$): The empty initial state ($S_{\emptyset}$) represents a completely unconfigured, bare operating system or a purely hypothetical state without applications and configurations. This state is used for theoretical purposes to define the starting point of a system’s configuration. $S_{\emptyset}\in\mathcal{S}$.
Base image ($B$): From the empty initial state ($S_{\emptyset}$), we can construct the base image $B$ by applying a series of actions. The base image $B$ is essentially the final state achieved by applying a sequence of instructions starting from $S_{\emptyset}$:
$B=\delta^{m}(S_{\emptyset},{C^{B}_{1},C^{B}_{2},\ldots,C^{B}_{m}})$ (2)
Here, ${C^{B}_{1},C^{B}_{2},\ldots,C^{B}_{m}}$ are the individual commands that constitute the base image.
2.3 Configuration process and target state
Configuration Process ($\Gamma$): The configuration process, denoted as $\Gamma$, involves transforming the environment starting from the base image $B$ by applying a series of actions. The result of this process is a state $S_{f}$, which we refer to as the resultant state. Formally, the configuration process can be represented as:
$S_{f}=\delta^{n}\left(B,{C_{1},C_{2},\ldots,C_{n}}\right)$ (3)
Following this transformation, it is crucial to verify whether $S_{f}$ meets all the required conditions to be considered the target state. This verification is performed using the state verification function.
State verification ($\epsilon$): The state verification function ($\epsilon$) determines whether the resultant state ($S_{f}$) successfully runs all tests within the repository after executing the configuration process ($\Gamma$). It maps a given state ($S$) to a binary value indicating whether all tests could be run in the resultant configuration: $\epsilon(S)$ returns 0 if all tests are successfully run, and 1 otherwise. If $\epsilon(S_{f})=0$, it indicates that the configuration process has finished and the resultant state $S_{f}$ is considered the target state. Otherwise, it means that $S_{f}$ does not meet the necessary conditions.
@@ 3 Repo2Run
Figure 3: An example process of Repo2Run, which illustrates two main parts: Event Stream and Repo2Run Workflow. The Event Stream tracks the entire action-observation process, where green boxes represent system prompts, yellow boxes represent responses from the LLM agent including the actions, and blue boxes represent observations from the Runtime Environment. The shaded actions indicate configurations abandoned after base image change. In the Event Stream, the blue text indicates that the command starts running, the green text indicates that the command runs successfully, and the red text indicates that the command fails. The Repo2Run Workflow consists of the internal environment and the external environment. The internal environment serves as the actual configuration Docker-based sandbox, which builds an actual testing runtime environment. The external environment executes the action-observation process and assists in the configuration process within the internal environment.
In this part, we introduce the design of Repo2Run and atomic configuration synthesis. As shown in Repo2Run Workflow in Figure 3, the dual-environment architecture consists of an external environment and an internal environment. The internal environment is a Docker container where the actual environment configuration is executed, while the external environment assists the internal environment in setting up the configuration. After successfully configuring the environment, the Dockerfile generator will create the Dockerfile.
Figure 4: An example of the Dockerfile generator to transfer the commands into a runnable Dockerfile.

### v3

# Repo2Run: Automated Building Executable Environment for Code Repository at Scale

@@ Abstract
Scaling up executable code data is significant for improving language models’ software engineering capability. The intricate nature of the process makes it labor-intensive, time-consuming and expert-knowledge-dependent to build a large number of executable code repositories, limiting the scalability of existing work based on running tests. The primary bottleneck lies in the automated building of test environments for different repositories, which is an essential yet underexplored task. To mitigate the gap, we introduce Repo2Run, the first LLM-based agent aiming at automating the building of executable test environments for any repositories at scale. Specifically, given a code repository, Repo2Run iteratively builds the Docker image, runs unit tests based on the feedback of the building, and synthesizes the Dockerfile until the entire pipeline is executed successfully. The resulting Dockerfile can then be used to create Docker container environments for running code and tests. We created a benchmark containing 420 Python repositories with unit tests for evaluation. The results illustrate that Repo2Run achieves an 86.0% success rate, outperforming SWE-agent by 77.0%. The resources of Repo2Run are available at https://github.com/bytedance/Repo2Run.
@@ 1 Introduction
Figure 1: The pipeline of code repository mining and manual environment building. Developers manually write Dockerfiles through iterative steps including base environment selection, dependency installation, test running, error handling, validating the environment by running unit tests.
Figure 2: Four error types (highlighted in different colors) that SWE-agent fails to resolve during environment building.
Figure 3: The illustration of a command executing failed and “polluting” the environment. (a) Failed commands like “commandA” can irreversibly “pollute” the environment by altering packages, files, or directories, making subsequent builds unstable. (b) To reproduce the changes, such a failed command “RUN commandA” needs to be added to the Dockerfile. However, adding it will lead to building failure.
Large language models (LLMs) have recently illustrated significant progress on solving software engineering issues [1], driving the advent of numerous coding LLM agents like MetaGPT [2], SWE-agent [3], OpenHands [4], Copilot [5], and Cursor [6]. However, the lack of training environment remains a significant challenge to advancing LLMs in software engineering — As shown in Figure 1, while static code is readily accessible through the GitHub API, building executable environments for code testing and patch validation demands extensive manual effort even for skilled developers  [^1] We conducted a study to show how industry developers struggle in building environments. See Appendix E.. This forces researchers to choose between (1) relying on unreliable static code metrics for rewarding (e.g., SWE-Fixer [7]) and SWE-RL [8], and (2) engaging developers to manually build limited repository environments (e.g., SWE-Gym [9] and SWE-Smith [10]) without scalability. Thus, an automated environment building infrastructure is urgently needed but remains underexplored.
In this paper, we introduce Repo2Run, an LLM-based agent designed to build executable environments. The key idea is that if the agent can successfully navigate the process of building a repository’s dependencies and running its unit tests, its actions can be recorded and replayed to programmatically synthesize a runnable Dockerfile. This enables automated creation of isolated and consistent environments across different platforms via only one line command “docker build”, eliminating “it works on my machine” problem [11].
Repo2Run addresses two key challenges in automating executable environment building:
  1. 1.
    Hard to explore a valid trajectory that can successfully build the environment. We found that without effective actions, the LLM agent would struggle to handle such a complex process. As shown in Figure 2, SWE-agent [3], a well-known LLM-based agent, often struggles with unresolved issues during environment building due to its lack of specialized actions. To address this challenge, Repo2Run uses an internal environment for building and an external environment for assistance, with nine actions designed to resolve these issues.
  2. 2.
    Failing to synthesize a runnable Dockerfile without execution failure. As shown in Figure 3, we observed that some incorrect commands in the trajectory can “pollute” the environment, leading to errors during synthesis. To tackle this challenge, we design an adaptive rollback action. When encountering failed commands like “commandA” that may cause “pollution” to the environment, Repo2Run restores the internal environment to its pre-execution state. This ensures the building process remains consistent and repeatable using the Dockerfile, avoiding unexpected errors.
To demonstrate Repo2Run’s effectiveness, we first created a benchmark of 420 latest Python repositories with unit tests from GitHub in 2024, as existing datasets include only a very small number of repositories [1]. We choose Python for two key reasons: (1) the most popular software engineering benchmarks (e.g., SWE-bench [1]) focus on Python repositories, and (2) Python is one of the most widely used programming languages today [12]. We then evaluated Repo2Run on this benchmark, finding it successfully built the environments for 361 repositories, achieving an 86.0% success rate, which is 63.9% higher than the leading baseline. Notably, Repo2Run outperforms the SWE-agent by 77.0%, demonstrating the effectiveness of the proposed agent.
We believe Repo2Run will serve as the foundational infrastructure, driving future research at the intersection of AI and Software Engineering and facilitating the community to efficiently scale up their executable code data and environments.
2 Formulation
In this part, we define the task of executable environment building. Given a repository $R$, executable environment building aims to determine a suitable base image $B$ and a building process $\mathcal{P}$ such that the resulting environment state $S$ satisfies the verification $\varepsilon$, as defined in Equation (1):
$\varepsilon(R,S)=0,\text{where}~{}S=\delta(B,\mathcal{P})$ (1)
Here, $\delta$ is the state transition function, and $\varepsilon(R,S)=0$ indicates that the verification $\varepsilon$ is satisfied.
@@ 2.1 State transition
$\bullet$ Environment state ($S$): The environment state $S\in\mathcal{S}$ represents the current state of the computer system, which encompasses all variables, files, cache, etc.
$\bullet$ Command sequence ($\mathcal{C}$): The command sequence $\mathcal{C}\in\mathbb{C}$ represents a set of individual commands. Each individual command $C\in\mathcal{C}$ refers to an instruction or action that can be executed in the environment via interfaces like bash, thereby changing the system state.
$\bullet$ State transition ($\delta$): The state transition $\delta$ is a function that defines the process through which the system transitions from the start state $S_{start}\in\mathcal{S}$ to the end state $S_{end}\in\mathcal{S}$ after executing of a command sequence $\mathcal{C}$, as defined in Equation (2):
$\delta:\mathcal{S}\times\mathbb{C}\rightarrow\mathcal{S},\quad\delta(S_{start},\mathcal{C})=S_{end}$ (2)
@@ 2.2 Base image
$\bullet$ Empty state ($S_{\emptyset}$): The empty state $S_{\emptyset}\in\mathcal{S}$ represents a completely bare operating system or a purely hypothetical state without any builds.
$\bullet$ Base image ($B$): The base image $B\in\mathcal{S}$ is a special type of environment state, typically managed by professional teams for user convenience, such as “python:3.10”. Starting from the empty state $S_{\emptyset}$, the base image $B$ can be created by executing a predefined command sequence $\mathcal{C}_{B}\in\mathbb{C}$, i.e., $B=\delta(S_{\emptyset},\mathcal{C}_{B})$. Users can utilize these base images by simply adding their names to the Dockerfiles.
2.3 Building process and state verification
$\bullet$ Building process ($\mathcal{P}$): The building process $\mathcal{P}\in\mathbb{C}$ is the command sequence designed to build the environment from the base image $B$. We denote the resulting state as $S_{f}\in\mathcal{S}$, where $S_{f}=\delta(B,\mathcal{P})$.
$\bullet$ State verification ($\varepsilon$): The state verification $\varepsilon$ is a Boolean function used to determine whether the resulting state $S_{f}$ successfully runs all tests in the repository $R$. $\varepsilon(R,S_{f})=0$ indicates that all tests in the repository $R$ can be successfully run. Otherwise, at least one test fails.
@@ 3 Repo2Run
Figure 4: The workflow of Repo2Run, involving two phases: the build phase and the record phase. The build phase utilizes a dual-environment architecture: the internal environment with five actions for environment building, while the external environment with three actions assists the internal environment. The record phase converts the validated command sequence into a runnable Dockerfile for reconstructing the executable environment. See Appendix A for more examples of these actions.
In this part, we introduce the design of Repo2Run. Given a code repository, Repo2Run synthesizes a runnable Dockerfile, enabling the building of an executable environment. As shown in Figure 4, this process is divided into two phases: the build phase and the record phase.
In build phase, Repo2Run employs a dual-environment architecture consisting of an internal environment and an external environment. The internal environment is a Docker container where the agent can execute actions restricted to a container, e.g., test running. The external environment is a terminal where the agent can execute actions prohibited within a container, e.g., base image change. Once all tests in the internal environment are successfully executed, Repo2Run transitions to the record phase.
In record phase, Repo2Run synthesizes a runnable Dockerfile based on the command sequence executed during the build phase. This Dockerfile serves as a precise record of the build process, and its execution can be viewed as a replay of the original build phase.
Next, we introduce the external environment, internal environment, and Dockerfile synthesizer.